Can someone help me solve this problem?
I need to subtract:

- 1/2 - (-5/8)
Thanks,
Sher

Answer by fractalier(2101)

Okay we have
- 1/2 - (-5/8) =
- 1/2 + 5/8 =
- 4/8 + 5/8 =
1/8